OpenStudy (zepp):

Find the values of the constants a and b for which the following function is continuous, but not differentiable. \[ f(n) = \left\{ \begin{array}{l l} ax+b, & \quad x>0;\\ \sin2x & \quad x \le 0.\\ \end{array} \right\} \]

OpenStudy (zepp):

Following the theorem of diffentiation; if a function is differentiable, it is continuous, so how am I going to do this?

OpenStudy (zepp):

@asnaseer Help please :D

OpenStudy (zepp):

@myininaya Help!

OpenStudy (zepp):

@radar

OpenStudy (radar):

The only help that I can offer is, I hope myininaya gets over here.

OpenStudy (zepp):

Ahh. D:

OpenStudy (asnaseer):

If ƒ is differentiable at a point x, then ƒ must also be continuous at x. see: en.wikipedia.org/wiki/Differentiable_function

OpenStudy (asnaseer):

are you sure you have stated the question correctly?

OpenStudy (asnaseer):

hmmm.... maybe the fact that f(x) = ax +b for x > 0 makes it non-differentiable at x=0?

OpenStudy (asnaseer):

the way I would tackle this in that case is as follows...

OpenStudy (asnaseer):

find f(0)

OpenStudy (asnaseer):

what do you get?

OpenStudy (zepp):

Hold on

OpenStudy (zepp):

b?

OpenStudy (asnaseer):

look carefully at the definition of f(x)

OpenStudy (asnaseer):

at x=0:\[f(x)=\sin(2x)\]

OpenStudy (asnaseer):

so f(0) = ?

OpenStudy (zepp):

0?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

now, next we look at the limit as x tends to \(0^+\)

OpenStudy (asnaseer):

what would that give?

OpenStudy (asnaseer):

i.e. what is the limit of f(x) as x tends to zero from the positive side of the x-axes?

OpenStudy (asnaseer):

remember on the positive x-axes, f(x) = ax + b

OpenStudy (zepp):

So we take the \[\large \lim_{ x\rightarrow 0 } \sin(2x)\]?

OpenStudy (asnaseer):

no - look at what I just typed ^^^

OpenStudy (zepp):

I'm confused. :|

OpenStudy (asnaseer):

what is the f(x) defined as for x>0?

OpenStudy (zepp):

ax+b?

OpenStudy (asnaseer):

correct, so we know on the positive side of the x-axes, f(x) = ax + b so what does this tend to as x tends to 0 (from the positive side)?

OpenStudy (asnaseer):

do you understand zepp?

OpenStudy (zepp):

x tends to 0+ is equal to x tends to 0-? Since the ax+b doesn't lie on the y-axis?

OpenStudy (asnaseer):

@arghya - please let zepp answer this

OpenStudy (asnaseer):

let me draw a diagram to help here

OpenStudy (zepp):